OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 73 and the last term is –101?

OpenStudy (anonymous):

a1 = 73 , nth term, an = -101 n = 30 by applying formula sum of arithmetic sequence = (n/2)(a1+an), sum = (30/2)*(73-101) = 15*-28 = -420

OpenStudy (anonymous):

Thank You (: